OpenStudy (anonymous):

If a hospital patient is given 100 milligrams of medicine which leaves the bloodstream at 14% per hour, how many milligrams of medicine will remain in the system after 10 hours? Use the function A(t) = Ie^rt.

OpenStudy (anonymous):

A(t) = 100e^(r*t) e^r = 1 - 14/100 = 86/100 r lne = ln(86/100) r = ln(86/100) r = -0.150823 A(t) = 100e^(-0.150823t) Let t = 10 A(10) = 100e^(-0.150823*10) = 22.1305 which is closest to 24.66. Hope this helps :)